angelbeasley18:

The formula for the area of a parallelogram is A = bh, where b is the base and h is the height. A parallelogram that is not drawn to scale. The height is labeled (x minus 4) centimeters. The base is labeled (2 x squared + 2 x minus 6) centimeters. Which simplified expression represents the area of the parallelogram? –4x3 + 14x – 24 square centimeters 2x3 – 6x2 – 14x + 24 square centimeters –4x3 – 14x + 24 square centimeters 2x3 + 6x2 + 14x + 24 square centimeters

jhonyy9:

so given that A = b*h and you know that the height h = x - 4 and the base b = 2x^2 +2x -6 and now just substitute these given expressions of h and b inside area formula in place of h and b and make the calcule of multiplication - in this way will get easy the right correct answer sure hope so much that i ve helped you understandably